Foods Galore is a major distributor to restaurants and other institutional food users.

Foods Galore buys cereal from a manufacturer for $23 per case. Annual demand for cereal is 225,000 cases, and the company believes that the demand is constant at 900 cases per day for each of the 250 days per year that it is open for business. Average lead time from the supplier for replenishment orders is eight days, and the company believes that it is also constant. The purchasing agent at Foods Galore believes that annual inventory carrying cost is 20 percent and that it costs $68 to prepare, send, and receive an order. Use Table 7-2.

a-1. How many cases of cereal should Foods Galore order each time it places an order? (Round your answer to the nearest whole number.)

economic quantity order:

a-2. What will be the average inventory? (Round your answer to 1 decimal place.)

average inventory:

a-3. What will be the inventory turnover rate? (Round your answer to 1 decimal place.)

inventory turnover:

a-4. Calculate the total annual cost based on a product cost of $23/unit. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

total annual cost:

Foods Galore conducts an in-depth analysis of its inventory management practices and discovers several flaws in its previous approach. First, they find that by ordering 17,000 or more cases each time, they can obtain a price of $21 per case from the supplier.

b-1. Calculate the total annual cost based on a product cost of $21/unit.

total annual cost:

b-2. What order quantity should Foods Galore place?

  

quantity to be ordered:

b-3. How much will they save annually by ordering the quantity shown in Part b-2 instead of Part a-1? (Round your answer to 2 decimal places.)

amount saved:

c. In its analysis, Foods Galore determined that demand and lead time are not constant. In fact, demand has a standard deviation of 67 cases per day and lead time has a standard deviation of 1.7 days. Foods Galore management wants to evaluate two service level policies. One policy would incur a 5 percent risk of stockout while waiting for replenishment, the other only a 1 percent risk of stockout. What would be the cost of carrying the safety stocks for each of the two policies? (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

5% risk out of stock:

1& risk out of stock:
Answer #1

Answer a:

Annual demand= demand per Month*12 225000
cost per order $                    68.00
item cost $                    23.00
holding cost percent 20.0%
holding cost per year= holding cost percent*item cost $                      4.60
Answer a.1 EOQ= SQRT(2*Cost per order*D/holding cost per year) 2579.2
#orders per year = Annual demand/ordered quantity 87.24
Answer a.2 Average inventory= EOQ/2 1289.59
annual holding cost= EOQ*holding cost per year/2 $                5,932.1
Total order cost= No. of orders*order cost $                5,932.1
total ordering+holding cost $            11,864.23
Purchasing cost= annual demand*item cost $            5,175,000
Answer a.4 Total cost= holding cost+order cost+purchase cost $      5,186,864.23

Answer a.3: Inventory turnover= Annual demand/Average inventory= 225,000/1289.59 = 174.5

Note: Chegg's guidelines mandate to solve one question (First question) at a time. Please post again for remaining questions.